# SAFE: Multitask Failure Detection for Vision-Language-Action Models

### v1

@@ Abstract
While vision-language-action models (VLAs) have shown promising robotic behaviors across a diverse set of manipulation tasks, they achieve limited success rates when deployed on novel tasks out-of-the-box. To allow these policies to safely interact with their environments, we need a failure detector that gives a timely alert such that the robot can stop, backtrack, or ask for help. However, existing failure detectors are trained and tested only on one or a few specific tasks, while VLAs require the detector to generalize and detect failures also in unseen tasks and novel environments. In this paper, we introduce the multitask failure detection problem and propose SAFE, a failure detector for generalist robot policies such as VLAs. We analyze the VLA feature space and find that VLAs have sufficient high-level knowledge about task success and failure, which is generic across different tasks. Based on this insight, we design SAFE to learn from VLA internal features and predict a single scalar indicating the likelihood of task failure. SAFE is trained on both successful and failed rollouts, and is evaluated on unseen tasks. SAFE is compatible with different policy architectures. We test it on OpenVLA, $\pi_{0}$, and $\pi_{0}$-FAST in both simulated and real-world environments extensively. We compare SAFE with diverse baselines and show that SAFE achieves state-of-the-art failure detection performance and the best trade-off between accuracy and detection time using conformal prediction. More qualitative results can be found at https://vla-safe.github.io/.
@@ 1 Introduction
Recently, scaling up robot manipulation datasets has enabled the development of large vision-language-action (VLA) models, which are generalist manipulation policies that can follow language instructions and accomplish a wide range of tasks [1, 2, 3, 4, 5, 6]. However, when VLAs are directly deployed on unseen tasks without collecting additional demonstrations and finetuning the model, they still suffer from limited success rates and a wide range of failure modes. This has been demonstrated by evaluations in recent work [2, 7, 4]: while their success rates on seen tasks are as high as 80-90%, those on unseen ones drop to 30-60% out-of-the-box. Therefore, to safely and reliably deploy VLA policies in the real world, it is important to promptly detect their potential failures.
@@
Most existing failure detection methods train a separate failure detector for each task, and evaluate the detector only on that task [8, 9, 10, 11, 12, 13, 14, 15, 16]. While these methods work well for specialist policies, they do not suit generalists like VLAs. VLAs are designed to accomplish diverse tasks and may frequently encounter novel task instructions and unseen environments during deployment. In such cases, it is impractical to exhaustively collect rollouts and train a failure detector for every new task. Some recent works introduce task-generic failure detectors, but they either require sampling multiple actions [17] or need to query a large VLM [18, 19], which poses significant inference overhead for large VLAs in the real world. This motivates the need for an efficient and multitask failure detector that can generalize to unseen tasks zero-shot and detect failures in a timely manner during the on-policy rollout of the VLA.
@@
In this paper, we focus on the multitask failure detection problem. This setting evaluates failure detection performance of a VLA policy without collecting rollouts or finetuning the failure detector on unseen tasks. To our knowledge, such multitask failure detection ability for VLAs has not been shown in the literature. To tackle this problem, we study the internal features of VLAs and find that they capture high-level knowledge about task success and failure. As shown in Fig. 1, such knowledge can effectively distinguish failed rollouts from successful ones and is generic across different tasks.
Figure 1: The VLA’s internal features capture high-level information about task success and failure. When the VLA is failing, even though from different tasks, the features fall in the same “failure zone”. This motivates SAFE, an efficient multitask failure detector that is based on VLA internal features and can generalize to unseen tasks. Plot (a) visualizes latent features of $\pi_{0}$-FAST on LIBERO-10 [20] using t-SNE [21]. For successful rollouts, their features are colored in blue; for failed rollouts, features are colored in the blue-red gradient based on timesteps. Plot (b) visualizes the same set of t-SNE features, colored by task ID. In (c), we show two example rollouts over time and mark their corresponding projected features in (a) and (b).
@@
Based on this insight, we introduce SAFE, a ScAlable Failure Estimation method that scales across diverse tasks for generalist policies like VLAs. SAFE takes in a VLA’s internal features and regresses them to a single scalar indicating the likelihood of failure. By training on successful and failed rollouts of multiple tasks, SAFE learns to identify task-generic representations for failure detection. To determine the threshold for failure detection, we adopt the functional conformal prediction (CP) [22, 8] framework and calibrate the prediction band on the seen tasks. We conduct failure detection experiments on OpenVLA [2], $\pi_{0}$ [4] and $\pi_{0}$-FAST [5], in both simulation and the real world. For evaluation, we adapt diverse baseline failure detection methods from both the LLM literature [23, 24] and the robot learning literature [8, 17] onto VLAs. SAFE and baselines are evaluated on both training tasks and a set of held-out tasks. Experiments show that SAFE outperforms other existing baselines and achieves the best trade-off between accuracy and timeliness for failure detection. The contributions of our paper can be summarized as follows:
@@
  • We analyze the VLA feature space and show that across different task instructions and environments, VLA’s internal features are separated for successful and failed rollouts.
@@
  • We propose SAFE, a multitask failure detector designed for generalist robot policies. By operating on latent features, training on multiple tasks, and using conformal prediction methods, SAFE shows generalization ability in detecting failures on unseen tasks.
@@ 2.1 Vision-Language-Action Models
Recent advances in large-scale machine learning and the availability of extensive robot demonstration datasets have paved the way for VLA models [1, 2, 3, 4, 7, 25, 26, 6]. These generalist robotic policies are initialized from pretrained large-scale VLMs [27, 28, 29], and thus inherit the strong ability to understand diverse semantic concepts from both images and language. They are augmented with an action head that produces continuous control signals, through per-step binning [1, 7, 2, 25], diffusion networks [3, 4, 30, 31] or frequency-space tokenization [5]. These VLAs are then trained on vast robotic datasets covering a wide array of tasks [32, 33, 34]. As a result, VLAs can successfully perform familiar tasks in new environments and even tackle previously unseen tasks when provided with novel language instructions. Nevertheless, significant variability in real-world deployments and the challenging domain gaps between training and testing environments continue to hinder VLA performance. Most state-of-the-art VLA models achieve success rates between 30% and 60% when evaluated out-of-the-box on real robots with unseen task instructions [2, 32, 4]. These limitations highlight the need for robust multitask failure detection methods tailored to generalist VLA models.
@@ 2.2 Failure Detection in Robot Manipulation
Monitoring failures is critical when deploying robotic policies in real-world environments, as even minor errors can result in hazardous conditions [35, 36, 37]. The literature on failure detection in robot learning can be broadly divided into unsupervised out-of-distribution (OOD) detection [8, 9, 10, 11] and supervised failure detection [9, 12, 13, 14, 15, 16]. OOD detection-based methods treat successful executions as the in-domain baseline and consider any deviation from this norm as a failure. However, the assumption that any unseen scenario constitutes a failure is overly restrictive for generalist VLAs, which may frequently encounter unseen tasks at test time. These unseen tasks are likely different from the in-domain training data but should not be simply treated as failures. Our proposed method, SAFE, falls within the supervised failure detection category, leveraging both successful and failed rollouts to train a failure classifier. Yet SAFE trains a single unified failure detector for all tasks, while other existing ones typically require training and calibration of separate classifiers for each task and have not yet demonstrated effectiveness on generalist policies like VLAs. Some recent works have explored the multitask failure detection task by designing action consistency scores [17] or instruction-finetuning a VLM [18, 19], but they require either sampling multiple actions or querying a large VLM, which poses significant overhead for controlling robots in real time.
Recently, FAIL-Detect [8] conducted a systematic evaluation of various failure detection methods, including OOD detection-based approaches [38, 39, 9], smoothness-based techniques [40], and consistency-based strategies [17]. Their experiments indicate that the best performance was achieved by LogpZO [38] and RND [39], both using a learned network to model data distribution and detect OOD states. However, their evaluation is limited to only single-task policies, and our evaluation in multitask setting shows that their best performing RND and LogpZO methods suffer from overfitting to the training tasks.
@@ 2.3 Uncertainty Quantification for LLM
Although LLMs and VLMs have demonstrated remarkable understanding and generative capabilities across various tasks, they are prone to producing hallucinated responses [41, 42, 43]. Numerous methods have been developed for uncertainty quantification (UQ) in LLMs/VLMs. Token-level uncertainty quantification methods estimate uncertainty by analyzing the probability distribution over each generated token to assess the likelihood of an entire response [44, 45, 46]. In contrast, semantic-similarity methods generate multiple responses to the same query and evaluate their semantic alignment [24, 47, 48]; a higher variance among responses typically signals low confidence. Since vision-language-action models (VLAs) share the generative nature and transformer architecture of LLMs/VLMs, we adapt these UQ methods to VLAs as promising baselines and evaluate their performance on failure detection. Recent research has also explored the internal latent space of LLMs for hallucination detection [49, 50, 51, 52, 53, 54, 55]. These methods train a classifier on internal latent features to distinguish between truthful and hallucinated outputs, paralleling supervised failure detection techniques in robotics. This approach has proven to be simple, efficient, and effective for UQ in LLMs. In our study, we investigate its application to large VLA policies and observe promising performance in robotic tasks.
@@ 3 Problem Formulation
This work aims to detect when a robot policy fails its task execution. Specifically, we aim for a multi-task failure detector that performs well when generalist VLAs perform novel tasks at inference time. At timestep $t$, a VLA is given an input observation $\mathbf{o}_{t}$, consisting of RGB images, natural language instruction, and current robot state, and outputs a control signal $\mathbf{A}_{t}=[\mathbf{a}_{t},\mathbf{a}_{t+1},\dots,\mathbf{a}_{t+H-1}]$, which is a chunk of actions for the next $H$ timesteps. The first $H^{\prime}$ ($H^{\prime}\leq H$) actions in $\mathbf{A}_{t}$ are executed, and then the VLA replans a new action sequence $\mathbf{A}_{t+H^{\prime}}$ at time $t+H^{\prime}$. We denote the internal embedding vector within the VLA model at time $t$ as $\mathbf{e}_{t}$. Some VLAs [1, 2, 4, 5] also decode a series of $m$ tokens $\mathbf{W}_{t}=[\mathbf{w}_{t}^{1},\dots,\mathbf{w}_{t}^{m}]$ before converting them into the actual action vector. To train and evaluate failure detection models, we run the VLA on different tasks in simulation or the real world, collect the rollout trajectory $\tau_{i}=\{(\mathbf{o}_{t},\mathbf{e}_{t},\mathbf{W}_{t},\mathbf{A}_{t})\}_{t=0,H^{\prime},\dots,nH^{\prime}}$ with time duration $T=nH^{\prime}$, and annotate each rollout with a failure label $y_{i}$ ($y_{i}=1$ if the robot fails to accomplish the task and $y_{i}=0$ if the robot succeeds). Note that for training, we only use the trajectory-level annotation $y_{i}$, and don’t require knowing the exact timestep when the policy starts to fail. A failure detector receives the rollout information up to time $t$ and predicts a failure score $s_{t}$, indicating the likelihood of task execution failure at time $t$. If $s_{t}$ exceeds a threshold $\delta_{t}$, a failure flag is raised, and then either the task execution is aborted or a human monitor will step in and take over the control. In this work, we use conformal prediction [56] to calibrate the threshold $\delta_{t}$.
@@ 4.1 Visual Analysis on VLA Latent Space
VLAs process multi-modal inputs and extract rich semantic information in their internal feature space. We hypothesize that these features also capture the high-level and abstract knowledge about task execution success/failure, by separating features from successful/failed rollouts into different regions. We study this hypothesis by visualizing the VLA features in Fig. 1, where we plot the internal features from $\pi_{0}$-FAST[5] when running the LIBERO-10 benchmark [20]. Fig. 1(a) demonstrates that when the VLA is failing, its internal features are grouped in the same region in the feature space (“failure zone”). Comparing Fig. 1(a) and Fig. 1(b), we can further see that although the features are extracted from different tasks with various instructions, objects and environments, when the VLA fails, its features fall in the same “failure zone”. Fig. 1(c) further illustrates how VLA’s features evolve in the feature space when VLA progresses temporally. From Fig. 1(c), we can see that failure rollout initially stays out of the “failure zone” when it progresses normally, and when the robot mistakenly drops the pot in the middle of execution and starts to fail, it steps into the “failure zone”. On the contrary, for the successful rollout, its features always stay out of the “failure zone”.
This visual analysis shows that the VLA’s internal features for succeeding and failing task executions are well separated in the feature space, and this separation is general across different tasks. Furthermore, during task execution, the features reflect how well the VLA performs on the current tasks in a timely manner. Inspired by this observation, we design SAFE, which uses the internal features of VLAs for failure detection.
Table 1: Failure detection results on simulation benchmarks, measured by area under ROC (ROC-AUC). “-” indicates that the failure detection method does not apply. Entries with gray background indicate the failure detection methods that sample 10 actions per inference timestep, while others use only 1 action. The first and second best-performing methods are colored in red and orange, respectively. Results are averaged over 3 random seeds with different splits of seen and unseen tasks.
[…]

### v2

@@ Abstract
While vision-language-action models (VLAs) have shown promising robotic behaviors across a diverse set of manipulation tasks, they achieve limited success rates when deployed on novel tasks out of the box. To allow these policies to safely interact with their environments, we need a failure detector that gives a timely alert such that the robot can stop, backtrack, or ask for help. However, existing failure detectors are trained and tested only on one or a few specific tasks, while generalist VLAs require the detector to generalize and detect failures also in unseen tasks and novel environments. In this paper, we introduce the multitask failure detection problem and propose SAFE, a failure detector for generalist robot policies such as VLAs. We analyze the VLA feature space and find that VLAs have sufficient high-level knowledge about task success and failure, which is generic across different tasks. Based on this insight, we design SAFE to learn from VLA internal features and predict a single scalar indicating the likelihood of task failure. SAFE is trained on both successful and failed rollouts, and is evaluated on unseen tasks. SAFE is compatible with different policy architectures. We test it on OpenVLA, $\pi_{0}$, and $\pi_{0}$-FAST in both simulated and real-world environments extensively. We compare SAFE with diverse baselines and show that SAFE achieves state-of-the-art failure detection performance and a favorable trade-off between accuracy and detection time using conformal prediction. More qualitative results and code can be found at the project webpage: https://vla-safe.github.io/.
@@ 1 Introduction
Recently, scaling up robot manipulation datasets has enabled the development of large vision-language-action (VLA) models, which are generalist manipulation policies that can follow language instructions and accomplish a wide range of tasks [1, 2, 3, 4, 5, 6]. However, when VLAs are directly deployed on unseen tasks without collecting additional demonstrations and finetuning the model, they still suffer from limited success rates and a wide range of failure modes. This has been demonstrated by evaluations in recent work [2, 7, 4]: while VLAs achieve success rates of 80–90% on seen tasks, their performance on unseen tasks drops to 30–60% out of the box. Therefore, to safely and reliably deploy VLA policies in the real world, it is important to promptly detect their potential failures.
@@
Most existing failure detection methods train a separate failure detector for each task, and evaluate the detector only on that task [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. While these methods work well for specialist policies, they do not suit generalists like VLAs. VLAs are designed to accomplish diverse tasks and may frequently encounter novel task instructions and unseen environments during deployment. In such cases, it is impractical to exhaustively collect rollouts and train a failure detector for every new task. Some recent works introduce task-generic failure detectors, but they either require sampling multiple actions [18] or need to query a large VLM [19, 20], which poses significant inference overhead for VLAs in the real world. This motivates the need for an efficient and multitask failure detector that can generalize to unseen tasks zero-shot and detect failures in a timely manner during the on-policy rollout of the VLA.
@@
In this paper, we focus on the multitask failure detection problem. This setting evaluates the failure detection performance of a VLA policy without collecting rollouts or finetuning the failure detector on unseen tasks. To our knowledge, such multitask failure detection capabilities for VLAs have not been shown in the literature. To tackle this problem, we study the internal features of VLAs and find that they capture high-level knowledge about task success and failure. As shown in Fig.˜1, failed rollouts occupy a distinct region (“failure zone”) in the VLA feature space, and this separation remains consistent across different tasks.
@@
Based on this insight, we introduce SAFE, a ScAlable Failure Estimation method that scales across diverse tasks for generalist policies like VLAs. SAFE takes in a VLA’s internal features and regresses them to a single scalar indicating the likelihood of failure. By training on successful and failed rollouts of multiple tasks, SAFE learns to identify task-generic representations for failure detection. To determine the threshold for failure detection, we adopt the functional conformal prediction (CP) [21, 8] framework and calibrate the prediction band on the seen tasks. We conduct failure detection experiments on OpenVLA [2], $\pi_{0}$ [4] and $\pi_{0}$-FAST [5], in both simulation and the real world. For evaluation, we adapt diverse baseline failure detection methods from both the LLM literature [22, 23] and the robot learning literature [8, 18] onto VLAs. SAFE and baselines are evaluated on both training (Seen) tasks and a set of held-out (Unseen) tasks. Experiments show that SAFE outperforms other existing baselines and achieves the best trade-off between accuracy and timeliness for failure detection. The contributions of our paper can be summarized as follows:
@@
  • We analyze the VLA feature space and show that, across different task instructions and environments, the internal features of the VLA distinctly separate successful and failed rollouts.
@@
  • We propose SAFE, a multitask failure detector designed for generalist robot policies. By operating on latent features, training on multiple tasks, and using conformal prediction methods, SAFE shows generalization capabilities in detecting failures on unseen tasks.
@@ 2.1 Vision-Language-Action Models
Recent advances in large-scale machine learning and the availability of extensive robot demonstration datasets have paved the way for VLA models [1, 2, 3, 4, 7, 24, 25, 6]. These generalist robotic policies are initialized from pretrained large-scale VLMs [26, 27, 28], and thus inherit the ability to understand diverse semantic concepts from both images and language. They are augmented with an action head that produces continuous control signals, through per-step binning [1, 7, 2, 24], diffusion networks [3, 4, 29, 30], or frequency-space tokenization [5]. These VLAs are then trained on vast robotic datasets covering a wide array of tasks [31, 32, 33]. As a result, VLAs can successfully perform familiar tasks in new environments and even tackle previously unseen tasks when provided with novel language instructions. Nevertheless, significant variability in real-world deployments and the challenging domain gaps between training and testing environments continue to hinder VLA performance. Most state-of-the-art VLA models achieve success rates between 30% and 60% when evaluated out-of-the-box on real robots with unseen task instructions [2, 31, 4]. These limitations highlight the need for robust multitask failure detection methods tailored to generalist VLA models.
@@ 2.2 Failure Detection in Robot Manipulation
Monitoring failures is critical when deploying robotic policies in real-world environments, as even minor errors can result in hazardous conditions [34, 35, 36]. The literature on failure detection in robot learning can be broadly divided into unsupervised out-of-distribution (OOD) detection [8, 9, 10, 11, 17] and supervised failure detection [9, 12, 13, 14, 15, 16, 37]. OOD detection-based methods treat successful executions as the in-domain baseline and consider any deviation from this norm as a failure. However, the assumption that any unseen scenario constitutes a failure is overly restrictive for generalist VLAs, which may frequently encounter unseen tasks at test time. These unseen tasks are likely different from the in-domain training data but should not be simply treated as failures. Our proposed method, SAFE, falls within the supervised failure detection category, leveraging both successful and failed rollouts to train a failure classifier. Unlike existing methods that train and calibrate separate classifiers per task, SAFE uses a single unified failure detector and works effectively on generalist policies like VLAs. Some recent works have explored multitask failure detection by designing action consistency scores [18] or instruction-finetuning a VLM [19, 20], but they require either sampling multiple actions or querying a large VLM, which poses significant overhead for controlling robots in real time.
Recently, FAIL-Detect [8] conducted a systematic evaluation of various failure detection methods, including OOD detection-based approaches [38, 39, 9], smoothness-based techniques [40], and consistency-based strategies [18]. Their experiments indicate that the best performance was achieved by LogpZO, which learns a proxy for the likelihood of the data in the observation embedding space using flow matching [8]. However, their evaluation is limited to only single-task policies, and our evaluation in the multitask setting shows that their best-performing LogpZO method suffers from overfitting to the training tasks.
@@ 2.3 Uncertainty Quantification for LLM
Although LLMs and VLMs have demonstrated remarkable understanding and generative capabilities across various tasks, they are prone to producing hallucinated responses [41, 42, 43]. Numerous methods have been developed for uncertainty quantification (UQ) in LLMs/VLMs. Token-level uncertainty quantification methods estimate uncertainty by analyzing the probability distribution over each generated token to assess the likelihood of an entire response [44, 45, 46]. Semantic-similarity methods generate multiple responses to the same query and evaluate their semantic alignment [23, 47, 48]; a higher variance among responses typically signals low confidence. Since vision-language-action models (VLAs) share the generative nature and transformer architecture of LLMs/VLMs, we adapt these UQ methods to VLAs as promising baselines and evaluate their performance on failure detection. Note that these UQ baselines are used as a proxy for failure detection, assuming that when a policy becomes uncertain about its actions, it will have a higher probability of failing the task. Recent research has also explored the internal latent space of LLMs for hallucination detection [49, 50, 51, 52, 53, 54, 55]. These methods train a classifier on internal latent features to distinguish between truthful and hallucinated outputs, paralleling supervised failure detection techniques in robotics. This approach has proven to be simple, efficient, and effective for UQ in LLMs. In our study, we investigate its application to large VLA policies and observe promising performance in robotic tasks.
Figure 1: The internal features of a VLA capture high-level information about task success and failure. When the VLA is failing, the features, even those from different tasks, fall into the same “failure zone”. This motivates SAFE, an efficient multitask failure detector that is based on VLA internal features and can generalize to unseen tasks. Plot (a) visualizes the latent features of $\pi_{0}$-FAST on LIBERO-10 [56] using t-SNE [57]. For successful rollouts, features are colored in blue. For failed rollouts, features follow a blue-to-red gradient based on timestep progression, with red corresponding to later timesteps that often coincide with failure. Plot (b) visualizes the same set of t-SNE features, colored by task ID. In (c), we show two example rollouts over time and mark their corresponding projected features in (a) and (b).
@@ 3 Problem Formulation
This work aims to detect when a robot policy fails during task execution. Specifically, we develop a multi-task failure detector that performs well when generalist VLAs encounter novel tasks at inference time. At timestep $t$, a VLA is given an input observation $\mathbf{o}_{t}$, consisting of RGB images, natural language instruction, and current robot state, and outputs a control signal $\mathbf{A}_{t}=[\mathbf{a}_{t},\mathbf{a}_{t+1},\dots,\mathbf{a}_{t+H-1}]$, which is a chunk of actions for the next $H$ timesteps. The first $H^{\prime}$ ($H^{\prime}\leq H$) actions in $\mathbf{A}_{t}$ are executed, and then the VLA replans a new action sequence $\mathbf{A}_{t+H^{\prime}}$ at time $t+H^{\prime}$. We denote the internal embedding vector within the VLA model at time $t$ as $\mathbf{e}_{t}$. Some VLAs [1, 2, 4, 5] also decode a series of $m$ tokens $\mathbf{W}_{t}=[\mathbf{w}_{t}^{1},\dots,\mathbf{w}_{t}^{m}]$ before converting them into the actual action vector. To train and evaluate failure detection models, we run the VLA on different tasks in simulation or the real world, collect the rollout trajectory $\tau_{i}=\{(\mathbf{o}_{t},\mathbf{e}_{t},\mathbf{W}_{t},\mathbf{A}_{t})\}_{t=0,H^{\prime},\dots,nH^{\prime}}$ with time duration $T=nH^{\prime}$, and annotate each rollout with a failure label $y_{i}$ ($y_{i}=1$ if the robot fails to accomplish the task and $y_{i}=0$ if the robot succeeds). Note that for training, we only use the trajectory-level annotation $y_{i}$, and do not require knowing the exact timestep when the policy starts to fail. A failure detector receives the rollout information up to time $t$ and predicts a failure score $s_{t}$, indicating the likelihood of task execution failure at time $t$. If $s_{t}$ exceeds a threshold $\delta_{t}$, a failure flag is raised, and then either the task execution is aborted or a human monitor will step in and take over the control. In this work, we use conformal prediction [58] to calibrate the threshold $\delta_{t}$.
@@ 4.1 Visual Analysis on VLA Latent Space
VLAs process multi-modal inputs and extract rich semantic information in their internal feature space. We hypothesize that these features also capture the high-level and abstract knowledge about task execution success/failure, by separating features from successful/failed rollouts into different regions. We study this hypothesis by visualizing the VLA features in Fig.˜1, where we plot the internal features from $\pi_{0}$-FAST[5] when running the LIBERO-10 benchmark [56]. Fig.˜1(a) demonstrates that when the VLA is failing, its internal features are grouped in the same region in the feature space (“failure zone”). Comparing Fig.˜1(a) and Fig.˜1(b), we can further see that although the features are extracted from different tasks with various instructions, objects and environments, when the VLA fails, its features fall in the same “failure zone”. Fig.˜1(c) further illustrates how VLA’s features evolve in the feature space when VLA progresses temporally. From Fig.˜1(c), we can see that failure rollout initially stays out of the “failure zone” when it progresses normally, and when the robot mistakenly drops the pot in the middle of execution and starts to fail, it steps into the “failure zone”. On the contrary, for the successful rollout, its features always stay out of the “failure zone”.
Figure 2: The proposed failure detector, SAFE, has three major components: (1) SAFE extracts the latent feature from the last layer of a VLA model; (2) SAFE sequentially processes the latent feature and predicts a failure score, using an MLP or an LSTM backbone; and (3) SAFE determines a time-varying threshold using functional conformal prediction (CP) on a hold-out calibration set. If the predicted score exceeds the threshold during testing, SAFE confidently detects a failure.
This visual analysis shows that the VLA’s internal features for succeeding and failing task executions are well separated in the feature space, and this separation is general across different tasks. Furthermore, during task execution, the features reflect how well the VLA performs on the current tasks in a timely manner. Inspired by this observation, we design SAFE, which uses the internal features of VLAs for failure detection. An overview of the proposed method is shown in Fig.˜2.